# GRAND: Graph Release with Assured Node Differential Privacy

### v2

@@ Appendix C Proof of Theorem 3 (Emprical CDF Convergence)
$\frac{1}{n}\sum_{i}\mathbf{1}_{\{|\tilde{Z}_{i2}-\bar{Z}_{i2}|\geq\varepsilon/\sqrt{2}\}}$ $\leq\frac{1}{n}\sum_{i}\mathbf{1}_{\{\tilde{Z}_{i1}\notin K\}}+\frac{1}{n}\sum_{i}\mathbf{1}_{\{\bar{Z}_{i1}\notin K\}}+\frac{1}{n}\sum_{i}\mathbf{1}_{\{|\tilde{Z}_{i1}-\bar{Z}_{i1}|\geq\delta\}}$
$+\frac{1}{n}\sum_{i}\mathbf{1}_{\{u_{i}\notin[a,b]\}}+\frac{1}{n}\sum_{i}\mathbf{1}_{\{v_{i}\notin[a,b]\}}+\frac{1}{n}\sum_{i}\mathbf{1}_{\{|u_{i}-v_{i}|\geq\delta_{1}\}}$
$\leq\gamma$
@@
Combining the above results and picking $\gamma=O(\varepsilon)$, we have
$\frac{1}{n}\sum_{i:\|\tilde{Z}_{i}-\bar{Z}_{i}\|\geq\varepsilon}1$ $\leq\frac{1}{n}\sum_{i}[\mathbf{1}_{\{|\tilde{Z}_{i1}-\bar{Z}_{i1}|\geq\varepsilon/\sqrt{2}\}}+\mathbf{1}_{\{|\tilde{Z}_{i2}-\bar{Z}_{i2}|\geq\varepsilon/\sqrt{2}\}}]=O_{P,m,n}(\varepsilon).$
@@ D.1 Consistency under inner product latent space models
In this section, we give the proof of the needed concentration bound of $\hat{Z}_{i}$’s. We will prove a high probability error bound for each individual $\hat{Z}_{i}$, $i=1,\cdots,n$. Recall that $\hat{Z}_{i}$’s are identifiable only up to an orthogonal transformation. This indeterminacy does not affect the results of our study, so we ignore it for notational simplicity.
The estimation of $\hat{Z}_{i}$, $i=1,\cdots,n$ is done by solving the problem
$\text{minimize}_{Z_{i},\alpha_{i}}~~\frac{1}{m}\sum_{j=n+1}^{m}-\big{[}A_{ij}\log(\sigma(Z_{i}^{T}\hat{Z}_{j}+\alpha_{i}+\alpha_{j}))+(1-A_{ij})\log(1-\sigma(Z_{i}^{T}\hat{Z}_{j}+\alpha_{i}+\alpha_{j})))\big{]}.$
@@
$\operatorname{P}(y_{i}=1\mid x_{i})=\sigma(\beta^{*\top}x_{i}),\quad\text{where }\sigma(z)=\frac{1}{1+e^{-z}}.$
@@
where $\ell(\beta;x,y)$ is the negative log-likelihood for a single data point (using the identity $\log(\sigma(z))=-\log(1+e^{-z})$ and $\log(1-\sigma(z))=-z-\log(1+e^{-z})$):
@@
For the proposed method, we observe perturbed predictors $\{\tilde{x}_{i}\}_{i=1}^{n}$ satisfying $\|{\tilde{x}_{i}-x_{i}}\|_{2}\leq\delta$. Following equation (D.1), we define two empirical average negative log-likelihood functions based on the sample $\{(x_{i},\tilde{x}_{i},y_{i})\}_{i=1}^{n}$:
@@
    $L(\beta)=\frac{1}{n}\sum_{i=1}^{n}\ell(\beta;\,x_{i},y_{i})=\frac{1}{n}\sum_{i=1}^{n}[\log(1+e^{\beta^{\top}x_{i}})-y_{i}(\beta^{\top}x_{i})].$
@@
    $\widetilde{L}(\beta)=\frac{1}{n}\sum_{i=1}^{n}\ell(\beta;\,\tilde{x}_{i},y_{i})=\frac{1}{n}\sum_{i=1}^{n}[\log(1+e^{\beta^{\top}\tilde{x}_{i}})-y_{i}(\beta^{\top}\tilde{x}_{i})].$
@@
  • The ”ideal” estimator (uncomputable in practice as $x_{i}$ are unknown):
    $\hat{\beta}^{\,(true)}=\operatorname{argmin}\displaylimits_{\beta\in\mathbb{R}^{d}}L(\beta).$
@@
    $\hat{\beta}=\operatorname{argmin}\displaylimits_{\beta\in\mathbb{R}^{d}}\widetilde{L}(\beta).$
Our goal is to bound the error $\|\hat{\beta}-\beta^{*}\|_{2}$ of the actual estimator relative to the true parameter $\beta^{*}$. Our theoretical analysis follows the strategy of Wang et al. [2024], with necessary modifications.
@@
    $x_{i}$ are i.i.d. sub-exponential vectors with bounded Orlicz norm and $E\|{x_{i}}\|^{4}<\infty$.
@@
    $\|\beta^{*}\|_{2}\leq R_{\beta}(n)=C_{\beta}\log m$ (constant $R_{\beta}>0$) for another (potentially large integer $m$).
@@
    $\mathcal{L}(\beta)$ minimized uniquely at $\beta^{*}$. Hessian $H^{*}:=\nabla^{2}\mathcal{L}(\beta^{*})\succeq\mu I$ ($\mu>0$). Exists fixed $r_{1}>0$ s.t. $\nabla^{2}\mathcal{L}(\beta)\succeq\mu/2I$ for $\beta\in B_{r_{1}}(\beta^{*})$.
@@
    $1/\sqrt{n}\ll\delta\ll 1$; $\delta=o\big{(}\frac{1}{\log n\log m}\big{)}$.
@@
Under the previous setup and conditions (C1)–(C5), for any desired polynomial decay rate $k>2$, there exists a sufficiently large constant $C_{1}$ and the corresponding $c_{1}$, such that the event $\mathcal{E}_{1}$: $\|\nabla L(\beta^{*})\|_{2}\leq C_{1}\sqrt{(\log n)/n}$ holds with probability at least $1-c_{1}n^{-(k+1)}$ for sufficiently large $n$.
@@
Let $g=\nabla L(\beta^{*})\in\mathbb{R}^{d}$. The $j$-th component is $g_{j}=\frac{1}{n}\sum_{i=1}^{n}Z_{ij}$, where $Z_{ij}=\epsilon_{i}x_{ij}=(\sigma(\beta^{*\top}x_{i})-y_{i})x_{ij}$. Since $x_{i}$ is sub-exponential and $\epsilon_{i}$ is bounded ($|\epsilon_{i}|\leq 1$), each $Z_{ij}$ is a mean-zero sub-exponential random variable. Let $K=\sup_{i,j}\|Z_{ij}\|_{\psi_{1}}$ be the sub-exponential parameter (Orlicz $\psi_{1}$-norm), which is $O(1)$. Let $\sigma^{2}=\sup_{i,j}\mathbb{E}[Z_{ij}^{2}]\leq\sup_{i,j}\frac{1}{4}\mathbb{E}[x_{ij}^{2}]=O(1)$.
@@
$\operatorname{P}\left(|g_{j}|\geq t\right)\leq 2\exp\left(-C_{B}n\min\left(\frac{t^{2}}{\sigma^{2}},\frac{t}{K}\right)\right)\leq 2\exp\left(-c^{\prime}n\min(t^{2},t)\right),$
for some constant $c^{\prime}$ depending on $\sigma^{2}$ and $K$. We want a bound $t_{n}$ such that $|g_{j}|\leq t_{n}$ holds for all $j=1,\dots,d$ simultaneously with probability $p_{n}^{\prime}\leq O(n^{-(k+1)})$ for the target $k>2$. Using a union bound:
$\operatorname{P}\left(\max_{j=1,\dots,d}|g_{j}|\geq t_{n}\right)\leq\sum_{j=1}^{d}\operatorname{P}\left(|g_{j}|\geq t_{n}\right)\leq d\cdot 2\exp\left(-c^{\prime}n\min(t_{n}^{2},t_{n})\right).$
Let $C_{1}^{\prime}=\sqrt{(k+1)/c^{\prime}}$. Choosing $t_{n}=C_{1}^{\prime}\sqrt{(\log n)/n}$ ensures the failure probability for a single coordinate is $O(n^{-(k+1)})$. This deviation $t_{n}\to 0$, justifying the use of $\min(t_{n}^{2},t_{n})=t_{n}^{2}$. By the union bound, the event $\mathcal{E}^{\prime}_{1}:=\{\max_{j}|g_{j}|\leq t_{n}\}$ holds with probability $P(\mathcal{E}^{\prime}_{1})\geq 1-O(n^{-(k+1)})$.
@@
$\|\nabla L(\beta^{*})\|_{2}^{2}=\sum_{j=1}^{d}g_{j}^{2}\leq\sum_{j=1}^{d}t_{n}^{2}=d\cdot t_{n}^{2}=d(C_{1}^{\prime})^{2}\frac{\log n}{n}.$
@@
$\|\nabla L(\beta^{*})\|_{2}\leq\sqrt{d}C_{1}^{\prime}\sqrt{\frac{\log n}{n}}.$
Let $C_{1}=\sqrt{d}C_{1}^{\prime}$. Thus, event $\mathcal{E}_{1}$ holds with probability $P(\mathcal{E}_{1})\geq 1-O(n^{-(k+1)})$.
@@
Under Assumptions (C1)–(C5), define event $\mathcal{E}_{2}$: $\nabla^{2}L(\beta)\succeq\frac{\mu}{4}I$ uniformly for $\beta\in B_{r_{1}}(\beta^{*})$. For some constant $c_{3}>0$, we have
$P(\mathcal{E}_{2})\geq 1-e^{-c_{3}n}$
@@
$X_{i}(\beta)\;=\;\nabla^{2}\ell(\beta;x_{i},y_{i})\;-\;\mathbb{E}\bigl{[}\nabla^{2}\ell(\beta;x_{i},y_{i})\bigr{]}.$
@@
$\nabla^{2}\ell(\beta;x,y)=\sigma^{\prime}(x^{\top}\beta)\,x\,x^{\top},\qquad\sigma^{\prime}(z)\;\leq\;\tfrac{1}{4},$
@@
$\|X_{i}(\beta)\|_{\rm op}\;\leq\;\bigl{\|}\nabla^{2}\ell(\beta;x_{i},y_{i})\bigr{\|}_{\rm op}+\bigl{\|}\mathbb{E}[\nabla^{2}\ell(\beta;x_{i},y_{i})]\bigr{\|}_{\rm op}\;=\;O(\|x_{i}\|^{2}).$
@@
$\sigma^{2}(\beta)=\Bigl{\|}\sum_{i=1}^{n}\mathbb{E}\bigl{[}X_{i}(\beta)^{2}\bigr{]}\Bigr{\|}_{\rm op}=O(n).$
@@
$\Pr\Bigl{(}\bigl{\|}\tfrac{1}{n}\sum_{i=1}^{n}X_{i}(\beta)\bigr{\|}_{\rm op}>t\Bigr{)}\;\leq\;2\,d\;\exp\!\Bigl{(}-c_{1}\,n\,\min\{t^{2}/K^{2},\;t/K\}\Bigr{)}.$
@@
$\Pr\Bigl{(}\|\nabla^{2}L(\beta)-\mathbb{E}\nabla^{2}\ell(\beta)\|_{\rm op}>\tfrac{\mu}{4}\Bigr{)}\;\leq\;2\,d\,\exp(-c_{2}\,n).$
@@
$B\;=\;\{\beta\in\mathbb{R}^{d}:\|\beta-\beta^{*}\|\leq r_{1}\}.$
@@
$\Pr\Bigl{(}\bigl{\|}\nabla^{2}L(\beta_{0})-\mathbb{E}\nabla^{2}\ell(\beta_{0})\bigr{\|}_{\rm op}>\tfrac{\mu}{4}\Bigr{)}\;\leq\;2\,d\,e^{-c_{2}n}.$
@@
$\Pr\bigl{(}\exists\,\beta_{0}\in\mathcal{N}:\|\nabla^{2}L(\beta_{0})-\mathbb{E}\nabla^{2}\ell(\beta_{0})\|_{\rm op}>\tfrac{\mu}{4}\bigr{)}\;\leq\;|\mathcal{N}|\;2\,d\,e^{-c_{2}n}\;=\;2\,d\,(3r_{1}/\varepsilon)^{d}\,e^{-c_{2}n}.$
@@
$\nabla^{2}\ell(\beta;x,y)-\nabla^{2}\ell(\beta^{\prime};x,y)=\bigl{[}\sigma^{\prime}(x^{\top}\beta)-\sigma^{\prime}(x^{\top}\beta^{\prime})\bigr{]}\,x\,x^{\top}.$
@@
$\nabla^{2}\ell(\beta;x,y)-\nabla^{2}\ell(\beta^{\prime};x,y)=\sigma^{\prime\prime}(\xi)\,\bigl{(}x^{\top}(\beta-\beta^{\prime})\bigr{)}\,x\,x^{\top}.$
@@
$\bigl{\|}\nabla^{2}\ell(\beta;x,y)-\nabla^{2}\ell(\beta^{\prime};x,y)\bigr{\|}_{\rm op}\;\leq\;\bigl{|}\sigma^{\prime\prime}(\xi)\bigr{|}\;\bigl{|}x^{\top}(\beta-\beta^{\prime})\bigr{|}\;\|x\|^{2}\;\leq\;\sup_{u}\bigl{|}\sigma^{\prime\prime}(u)\bigr{|}\;\|x\|\;\|\beta-\beta^{\prime}\|\;\|x\|^{2}.$
@@
$L\;=\;\Bigl{(}\sup_{u}|\sigma^{\prime\prime}(u)|\Bigr{)}\,\|x\|^{3}\;=\;O\bigl{(}\|x\|^{3}\bigr{)},$
@@
With probability at least $1-e^{-c_{3}^{\prime}n}$, we have $\max_{i\leq n}\|x_{i}\|\leq C\log n$. We can take
@@
$\varepsilon=\min\!\Bigl{\{}\frac{\mu}{16\,L},\,r_{1}\Bigr{\}}.$
@@
$\|\nabla^{2}L(\beta)-\nabla^{2}L(\beta_{0})\|_{\rm op}\;\leq\;L\,\|\beta-\beta_{0}\|\;\leq\;\frac{\mu}{16}.$
Also note that with this choice of $\varepsilon$, the probability in (i) holds with $O(e^{-c_{3}n})$.
On the intersection of the high‐probability events from (i) and the bound $\max_{i}\|x_{i}\|\leq C\log n$, we have for every $\beta\in B$:
$\|\nabla^{2}L(\beta)-\mathbb{E}\nabla^{2}\ell(\beta)\|_{\rm op}$ $\leq\|\nabla^{2}L(\beta_{0})-\mathbb{E}\nabla^{2}\ell(\beta_{0})\|_{\rm op}+\|\nabla^{2}L(\beta)-\nabla^{2}L(\beta_{0})\|_{\rm op}$
@@
Since $\mathbb{E}\nabla^{2}\ell(\beta)=\nabla^{2}\mathcal{L}(\beta)\succeq\mu I$, it follows that $\nabla^{2}L(\beta)\succeq\mu I-\tfrac{\mu}{2}I=\tfrac{\mu}{2}I>\tfrac{\mu}{4}I$ uniformly over $B$. Hence
$\Pr(\mathcal{E}_{2}^{c})\;=\;O(e^{-c_{3}n}).$
@@
$\operatorname{P}\left(\|\hat{\beta}-\beta^{*}\|_{2}\leq C\max\left\{\delta(\log n)(\log m),\frac{\log n}{\sqrt{n}}\right\}\right)\geq 1-C_{p}n^{-(k+1)}.$

### v3

@@ Appendix C Proof of Theorem 3 (Emprical CDF Convergence)
$\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}_{\{|\tilde{Z}_{i2}-\bar{Z}_{i2}|\geq\varepsilon/\sqrt{2}\}}$ $\leq\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}_{\{\tilde{Z}_{i1}\notin K\}}+\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}_{\{\bar{Z}_{i1}\notin K\}}+\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}_{\{|\tilde{Z}_{i1}-\bar{Z}_{i1}|\geq\delta\}}$
$\quad+\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}_{\{u_{i}\notin[a,b]\}}+\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}_{\{v_{i}\notin[a,b]\}}+\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}_{\{|u_{i}-v_{i}|\geq\delta_{1}\}}$
$\leq\gamma.$
@@
Combining the above results and picking $\gamma=\mathcal{O}(\varepsilon)$, we have
$\frac{1}{n}\sum_{i:\|\tilde{Z}_{i}-\bar{Z}_{i}\|\geq\varepsilon}1$ $\leq\frac{1}{n}\sum_{i=1}^{n}\left[\mathbf{1}_{\{|\tilde{Z}_{i1}-\bar{Z}_{i1}|\geq\varepsilon/\sqrt{2}\}}+\mathbf{1}_{\{|\tilde{Z}_{i2}-\bar{Z}_{i2}|\geq\varepsilon/\sqrt{2}\}}\right]=\mathcal{O}_{\mathbb{P},m,n}(\varepsilon).$
@@ D.1 Consistency under inner product latent space models
In this section, we give the proof of the needed concentration bound of $\hat{Z}_{i}$’s. We will prove a high probability error bound for each individual $\hat{Z}_{i}$, $i=1,\ldots,n$. Recall that $\hat{Z}_{i}$’s are identifiable only up to an orthogonal transformation. This indeterminacy does not affect the results of our study, so we ignore it for notational simplicity.
The estimation of $\hat{Z}_{i}$, $i=1,\ldots,n$ is done by solving the problem
$\mathop{\arg\min}_{(X_{i},\alpha_{i})\in\mathbb{R}^{d}}\sum_{j=n+1}^{n+m}-\{A_{ij}\log(\sigma(X_{i}^{\top}\hat{X}_{j}+\alpha_{i}+\hat{\alpha}_{j}))+(1-A_{ij})\log(1-\sigma(X_{i}^{\top}\hat{X}_{j}+\alpha_{i}+\hat{\alpha}_{j}))\}.$
@@
$\mathbb{P}(y_{i}=1\,|\,x_{i})=\sigma(\beta^{*\top}x_{i}),\quad\text{where }\sigma(z)=\frac{1}{1+e^{-z}}.$
@@
where $\ell(\beta;x,y)$ is the negative log-likelihood for a single data point (using the identity $\log(\sigma(z))=z-\log(1+e^{z})$ and $\log(1-\sigma(z))=-\log(1+e^{z})$):
@@
For the proposed method, we observe perturbed predictors $\{\tilde{x}_{i}\}_{i=1}^{n}$ satisfying $\|{\tilde{x}_{i}-x_{i}}\|\leq\delta$. Following equation (D.1), we define two empirical average negative log-likelihood functions based on the sample $\{(x_{i},\tilde{x}_{i},y_{i})\}_{i=1}^{n}$:
@@
    $\mathcal{L}(\beta)=\frac{1}{n}\sum_{i=1}^{n}\ell(\beta;x_{i},y_{i})=\frac{1}{n}\sum_{i=1}^{n}[\log(1+e^{\beta^{\top}x_{i}})-y_{i}(\beta^{\top}x_{i})].$
@@
    $\tilde{\mathcal{L}}(\beta)=\frac{1}{n}\sum_{i=1}^{n}\ell(\beta;\tilde{x}_{i},y_{i})=\frac{1}{n}\sum_{i=1}^{n}[\log(1+e^{\beta^{\top}\tilde{x}_{i}})-y_{i}(\beta^{\top}\tilde{x}_{i})].$
@@
  • The ideal estimator (uncomputable in practice as $x_{i}$ are unknown):
    $\hat{\beta}_{\rm true}=\mathop{\arg\min}_{\beta\in\mathbb{R}^{d}}\mathcal{L}(\beta).$
@@
    $\hat{\beta}=\mathop{\arg\min}_{\beta\in\mathbb{R}^{d}}\tilde{\mathcal{L}}(\beta).$
Our goal is to bound the error $\|\hat{\beta}-\beta^{*}\|$ of the actual estimator relative to the true parameter $\beta^{*}$. Our theoretical analysis follows the strategy of Wang et al. [2024], with necessary modifications.
@@
    $x_{i}$ are i.i.d. sub-exponential vectors with bounded Orlicz norm and $\mathbb{E}\|{x_{i}}\|^{4}<\infty$.
@@
    $\|\beta^{*}\|\leq R_{\beta}(n)=C_{\beta}\log m$ (constant $R_{\beta}>0$) for another (potentially large integer $m$).
@@
    $\mathcal{L}(\beta)$ minimized uniquely at $\beta^{*}$. Hessian $H^{*}:=\nabla^{2}\mathcal{L}(\beta^{*})\succeq\mu I_{d}$ ($\mu>0$). Exists fixed $r_{1}>0$ s.t. $\nabla^{2}\mathcal{L}(\beta)\succeq(\mu/2)I_{d}$ for $\beta\in B_{r_{1}}(\beta^{*})$.
@@
    $1/\sqrt{n}\ll\delta\ll 1$; $\delta=o(1/(\log n\log m))$.
@@
Under the previous setup and Assumptions (C1)–(C5), for any desired polynomial decay rate $k>2$, there exists a sufficiently large constant $C_{1}$ and the corresponding $c_{1}$, such that the event $\mathcal{E}_{1}$: $\|\nabla\mathcal{L}(\beta^{*})\|\leq C_{1}\sqrt{(\log n)/n}$ holds with probability at least $1-c_{1}n^{-(k+1)}$ for sufficiently large $n$.
@@
Let $g=\nabla\mathcal{L}(\beta^{*})\in\mathbb{R}^{d}$. The $j$-th component is $g_{j}=\frac{1}{n}\sum_{i=1}^{n}Z_{ij}$, where $Z_{ij}=\varepsilon_{i}x_{ij}=(\sigma(\beta^{*\top}x_{i})-y_{i})x_{ij}$. Since $x_{i}$ is sub-exponential and $\varepsilon_{i}$ is bounded ($|\varepsilon_{i}|\leq 1$), each $Z_{ij}$ is a mean-zero sub-exponential random variable. Let $K=\sup_{i,j}\|Z_{ij}\|_{\psi_{1}}$ be the sub-exponential parameter (Orlicz $\psi_{1}$-norm), which is $\mathcal{O}(1)$. Let $\sigma^{2}=\sup_{i,j}\mathbb{E}[Z_{ij}^{2}]\leq\sup_{i,j}\frac{1}{4}\mathbb{E}[x_{ij}^{2}]=\mathcal{O}(1)$.
@@
$\mathbb{P}(|g_{j}|\geq t)\leq 2\exp\left(-C_{B}n\min\left\{\frac{t^{2}}{\sigma^{2}},\frac{t}{K}\right\}\right)\leq 2\exp(-c^{\prime}n\min\{t^{2},t\})$
for some constant $c^{\prime}$ depending on $\sigma^{2}$ and $K$. We want a bound $t_{n}$ such that $|g_{j}|\leq t_{n}$ holds for all $j=1,\dots,d$ simultaneously with probability $p_{n}^{\prime}\leq\mathcal{O}(n^{-(k+1)})$ for the target $k>2$. Using a union bound:
$\mathbb{P}\left(\max_{1\leq j\leq d}|g_{j}|\geq t_{n}\right)\leq\sum_{j=1}^{d}\mathbb{P}(|g_{j}|\geq t_{n})\leq d\cdot 2\exp(-c^{\prime}n\min\{t_{n}^{2},t_{n}\}).$
Let $C_{1}^{\prime}=\sqrt{(k+1)/c^{\prime}}$. Choosing $t_{n}=C_{1}^{\prime}\sqrt{(\log n)/n}$ ensures the failure probability for a single coordinate is $\mathcal{O}(n^{-(k+1)})$. This deviation $t_{n}\to 0$, justifying the use of $\min\{t_{n}^{2},t_{n}\}=t_{n}^{2}$. By the union bound, the event $\mathcal{E}^{\prime}_{1}:=\left\{\max_{j\in[d]}|g_{j}|\leq t_{n}\right\}$ holds with probability $\mathbb{P}(\mathcal{E}^{\prime}_{1})\geq 1-\mathcal{O}(n^{-(k+1)})$.
@@
$\|\nabla\mathcal{L}(\beta^{*})\|^{2}=\sum_{j=1}^{d}g_{j}^{2}\leq\sum_{j=1}^{d}t_{n}^{2}=d\cdot t_{n}^{2}=d(C_{1}^{\prime})^{2}\frac{\log n}{n}.$
@@
$\|\nabla\mathcal{L}(\beta^{*})\|\leq\sqrt{d}C_{1}^{\prime}\sqrt{\frac{\log n}{n}}.$
Let $C_{1}=\sqrt{d}C_{1}^{\prime}$. Thus, event $\mathcal{E}_{1}$ holds with probability $\mathbb{P}(\mathcal{E}_{1})\geq 1-\mathcal{O}(n^{-(k+1)})$.
@@
Under Assumptions (C1)–(C5), define event $\mathcal{E}_{2}$: $\nabla^{2}\mathcal{L}(\beta)\succeq(\mu/4)I_{d}$ uniformly for $\beta\in B_{r_{1}}(\beta^{*})$. For some constant $c_{3}>0$, we have
$\mathbb{P}(\mathcal{E}_{2})\geq 1-e^{-c_{3}n}$
@@
$X_{i}(\beta)=\nabla^{2}\ell(\beta;x_{i},y_{i})-\mathbb{E}[\nabla^{2}\ell(\beta;x_{i},y_{i})].$
@@
$\nabla^{2}\ell(\beta;x,y)=\sigma^{\prime}(x^{\top}\beta)xx^{\top},\quad\sigma^{\prime}(z)\leq\frac{1}{4},$
@@
$\|X_{i}(\beta)\|_{\rm op}\leq\|\nabla^{2}\ell(\beta;x_{i},y_{i})\|_{\rm op}+\|\mathbb{E}[\nabla^{2}\ell(\beta;x_{i},y_{i})]\|_{\rm op}=\mathcal{O}(\|x_{i}\|^{2}).$
@@
$\sigma^{2}(\beta)=\left\|\sum_{i=1}^{n}\mathbb{E}[X_{i}(\beta)^{2}]\right\|_{\rm op}=\mathcal{O}(n).$
@@
$\mathbb{P}\left(\left\|\frac{1}{n}\sum_{i=1}^{n}X_{i}(\beta)\right\|_{\rm op}>t\right)\leq 2d\exp\left(-c_{1}n\min\Bigl\{\frac{t^{2}}{K^{2}},\frac{t}{K}\Bigr\}\right).$
@@
$\mathbb{P}\left(\|\nabla^{2}\mathcal{L}(\beta)-\mathbb{E}[\nabla^{2}\ell(\beta)]\|_{\rm op}>\frac{\mu}{4}\right)\leq 2de^{-c_{2}n}.$
@@
$B=\left\{\beta\in\mathbb{R}^{d}:\|\beta-\beta^{*}\|\leq r_{1}\right\}.$
@@
$\mathbb{P}\left(\|\nabla^{2}\mathcal{L}(\beta_{0})-\mathbb{E}[\nabla^{2}\ell(\beta_{0})]\|_{\rm op}>\frac{\mu}{4}\right)\leq 2de^{-c_{2}n}.$
@@
$\mathbb{P}\left(\exists\beta_{0}\in\mathcal{N}:\|\nabla^{2}\mathcal{L}(\beta_{0})-\mathbb{E}[\nabla^{2}\ell(\beta_{0})]\|_{\rm op}>\frac{\mu}{4}\right)\leq|\mathcal{N}|\cdot 2de^{-c_{2}n}=2d\Bigl(\frac{3r_{1}}{\varepsilon}\Bigr)^{d}e^{-c_{2}n}.$
@@
$\nabla^{2}\ell(\beta;x,y)-\nabla^{2}\ell(\beta^{\prime};x,y)=[\sigma^{\prime}(x^{\top}\beta)-\sigma^{\prime}(x^{\top}\beta^{\prime})]xx^{\top}.$
@@
$\nabla^{2}\ell(\beta;x,y)-\nabla^{2}\ell(\beta^{\prime};x,y)=\sigma^{\prime\prime}(\xi)[x^{\top}(\beta-\beta^{\prime})]xx^{\top}.$
@@
$\|\nabla^{2}\ell(\beta;x,y)-\nabla^{2}\ell(\beta^{\prime};x,y)\|_{\rm op}\leq|\sigma^{\prime\prime}(\xi)||x^{\top}(\beta-\beta^{\prime})|\|x\|^{2}\leq\sup_{u}\bigl|\sigma^{\prime\prime}(u)\bigr|\|x\|\|\beta-\beta^{\prime}\|\|x\|^{2}.$
@@
$L=\Bigl(\sup_{u}|\sigma^{\prime\prime}(u)|\Bigr)\|x\|^{3}=\mathcal{O}(\|x\|^{3}),$
@@
With probability at least $1-e^{-c_{3}^{\prime}n}$, we have $\max_{i\in[n]}\|x_{i}\|\leq C\log n$. We can take
@@
$\varepsilon=\min\left\{\frac{\mu}{16L},r_{1}\right\}.$
@@
$\|\nabla^{2}\mathcal{L}(\beta)-\nabla^{2}\mathcal{L}(\beta_{0})\|_{\rm op}\leq L\|\beta-\beta_{0}\|\leq\frac{\mu}{16}.$
Also note that with this choice of $\varepsilon$, the probability in (i) holds with $\mathcal{O}(e^{-c_{3}n})$.
On the intersection of the high‐probability events from (i) and the bound $\max_{i\in[n]}\|x_{i}\|\leq C\log n$, we have for every $\beta\in B$:
$\|\nabla^{2}\mathcal{L}(\beta)-\mathbb{E}[\nabla^{2}\ell(\beta)]\|_{\rm op}$ $\leq\|\nabla^{2}\mathcal{L}(\beta_{0})-\mathbb{E}[\nabla^{2}\ell(\beta_{0})]\|_{\rm op}+\|\nabla^{2}\mathcal{L}(\beta)-\nabla^{2}\mathcal{L}(\beta_{0})\|_{\rm op}$
@@
Since $\mathbb{E}[\nabla^{2}\ell(\beta)]=\nabla^{2}\mathcal{L}(\beta)\succeq\mu I_{d}$, it follows that $\nabla^{2}L(\beta)\succeq\mu I_{d}-(\mu/2)I_{d}=(\mu/2)I_{d}\succ(\mu/4)I_{d}$ uniformly over $B$. Hence
$\mathbb{P}(\mathcal{E}_{2}^{c})=\mathcal{O}(e^{-c_{3}n}).$
@@
$\mathbb{P}\left(\|\hat{\beta}-\beta^{*}\|\leq C\max\left\{\delta\log n\log m,\frac{\log n}{\sqrt{n}}\right\}\right)\geq 1-C_{p}n^{-(k+1)}.$